OpenStudy (anonymous):

c=11.4, a=10.2. find b

OpenStudy (anonymous):

Are these sides of a right triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

If that's a right triangle, then a^2 + b^2 = c^2

OpenStudy (anonymous):

\[\sqrt{11.4 ^2 -10.2^2}=\sqrt{129.96-104.04}=\sqrt{25.92}≈5.1\]